# Construction of Optimal Algorithms for Function Approximation in Gaussian Sobolev Spaces

### v2

@@ 1 Introduction
for some $\alpha\in\mathbb{N}$ and $p<q<\infty$. This space consists of functions whose weak derivatives up to order $\alpha$ are in $L_{\rho}^{q}$.
Function approximation in high dimension is an important task. The sampling recovery problem in Gaussian Sobolev spaces has been recently studied by a number of authors LPE2023; DK2023; GHRR2023. In this paper, we study sampling recovery in the one-dimensional Gaussian Sobolev spaces $W_{\rho}^{\alpha,q}$ and construct explicit linear algorithms that attain the optimal or almost optimal rate of worst-case convergence. It has been shown by Dũng and Nguyen (DK2023, Theorem 3.3) that the best possible convergence rate of sampling recovery in $W_{\rho}^{\alpha,q}$ for $p<q=2$ is of order $n^{-\alpha}$, and we show two alternative methods achieving this rate with explicit computational cost. As one-dimensional algorithms often serve as a foundation for their high-dimensional counterparts, our results motivate further study of such optimal algorithms in a high-dimensional setting.
@@
Numerical integration is closely related to function approximation. Here we mention results related to numerical integration over Gaussian Sobolev spaces for $q=2$. For the deterministic worst-case error, it has been shown that the rate $n^{-\alpha}$ is optimal by Dick et al. (DILP2018, Theorem 1), meaning that any deterministic linear quadrature cannot achieve a convergence rate faster than $n^{-\alpha}$. We note that their result is shown for so-called Hermite spaces, but the norm equivalence between Hermite spaces and Gaussian Sobolev spaces with $q=2$ is shown in (DILP2018, Lemma 6) and (KSG2022, Lemma 2.1). Recently, Dũng and Nguyen (DK2023, Theorem 2.3) showed that this rate is unchanged for general $q\in(1,\infty)$. The measure being Gaussian, Gauss–Hermite quadrature is a natural choice. However, Kazashi, Goda and one of the present authors KSG2022 have shown that Gauss–Hermite quadrature achieves the deterministic worst-case rate of only $n^{-\alpha/2}$. An upper bound of order $n^{-\alpha/2}$ was essentially obtained already by Mastroianni and Monegato MM1994 and the matching lower bound was proved in (KSG2022, Theorem 3.2). In contrast, a suitably truncated trapezoidal rule is shown to achieve the optimal rate $n^{-\alpha}$ up to a logarithmic factor in (KSG2022, Theorem 4.5). The subsequent work of these authors GKS2023 considers the randomized setting, i.e., algorithms are allowed to be random and their quality is measured in the worst-case root-mean-squared error (RMSE). Therein it is proved that no nonlinear adaptive algorithm can converge with a worst-case RMSE rate faster than $n^{-\alpha-1/2}$, and a randomized trapezoidal rule is shown to attain this rate. The strategy used for obtaining upper bounds of trapezoidal rules both in KSG2022; GKS2023 is based on introducing an auxiliary periodic function. This strategy is first used in NS2023, and we adapt this strategy for function approximation in Section 2.
Upon completing this research in the paper, we learned the paper D2025 by Dũng. The methods presented therein achieve the optimal rate of convergence, however they depend on various indices. We emphasize that our methods presented in this paper offer computationally cheap alternatives, especially the trigonometric interpolation.
@@ 2 Trigonometric interpolation
This section introduces our first algorithm based on trigonometric interpolation and analyses its $L^{p}_{\rho}$-error in terms of the number of function evaluation $n$. The algorithm is denoted by $A^{\dagger}_{n}:W^{\alpha,q}_{\rho}\to L^{p}_{\rho}$.
@@
First we note that $1/q-1/p+\varepsilon<0$. In the following, we use the Sobolev inequality for $F_{\tau}(x)\coloneqq\rho^{1/q-1/p+\varepsilon}(x)g^{(\tau)}(x)$. That is, from the boundedness of $\|F_{\tau}\|_{L^{q}(\mathbb{R})}$ and $\|F_{\tau}^{\prime}\|_{L^{q}(\mathbb{R})}$ we deduce the boundedness of $\|F_{\tau}\|_{L^{\infty}(\mathbb{R})}$. Let $H_{\ell}(x)$ be the $\ell$-th degree probabilists’ Hermite polynomial,
@@
First notice from the definition of Hermite polynomial that
@@
where in the last line, again we used $\|f^{(\tau-\ell+1)}(x)\|_{L^{q}_{\rho}}\leq\|f\|_{W^{\alpha,q}_{\rho}}<\infty$. Thus we have $\|g\|_{\textup{decay}}<\infty$. ∎
@@
The proof follows the strategy of (KSG2022, Lemma 4.1), where the case with $p=1$ and $q=2$ was proved. Recall that for $\tau,\tau^{\prime}\in\mathbb{N}_{0}$ scaled Bernoulli polynomials satisfy
@@
Next, we note that $G^{(\alpha)}$ and $g^{(\alpha)}$ only differ by the constant $\frac{1}{2T}\int_{-T}^{T}g^{(\alpha)}(y)\,\mathrm{d}y$ on [T,T][-T,T]. Since $f^{(\alpha)}\in L^{q}_{\rho}$, we know $g^{(\alpha)}$ is in $L^{q}([-T,T])$. Due to matching boundary values $G^{(\alpha-1)}(-T)=G^{(\alpha-1)}(T)$ and absolute continuity of $G^{(\alpha-1)}$ on $(-T-\delta,T+\delta)$, we have $\|G^{(\alpha)}\|_{L^{p}(\mathbb{T}([-T,T]))}<\infty$. ∎
@@
where we used the facts that $H_{\ell}(\cdot/\sqrt{p})\rho^{(q-p-\varepsilon)/p}$ is bounded on the real line and that $\rho(x)$ is decreasing in $\lvert x\rvert$. ∎
@@
The claim then follows from Lemma 2.4 with a=Ta=-T, $b=T$, and $\varepsilon=0$. ∎
@@
The result by Temlyakov (T2018_book, Theorem 2.4.4), where the interpolation operator is denoted by $I_{m}$, applies to our setting. Therein, the function space considered is defined via convolution kernels, but the norm equivalence to periodic Sobolev spaces
@@
is shown in (U2007, Theorem 2.7). Hence, by letting $S_{T}(x)\coloneqq Tx/\pi-T$, (T2018_book, Theorem 2.4.4) leads to
@@
Thus we have proved the claim. ∎
@@
where in the penultimate line, we used $|\frac{B^{[-T,T]}_{\tau}(x)}{\tau!}|\leq\frac{(2T)^{\tau-1}}{2}$ for $x\in[-T,T]$; see (NS2023, Equation (6)) or L1940.
@@
where this explicit constant here can be found, e.g., in (SN2016, Theorem 2.1), or we refer to EZ1966 as an earlier result.
@@
Hence we proved the claim. ∎
@@
This proves the claim.∎
@@
$\|f-A^{\dagger}_{n}(f)\|_{L_{\rho}^{p}}\leq C_{p,q,\alpha,\varepsilon}\|f\|_{W^{\alpha,q}_{\rho}}(\ln n)^{\frac{\alpha}{2}+\frac{1}{2}+\frac{1}{2p}}n^{-\alpha}.$
@@
This proof strategy of introducing the auxiliary function $G$ can be applied to other algorithms or problems. Let us elaborate this part here. Consider an abstract problem of approximating a linear operator or a linear functional $S:\mathcal{H}_{1}\to\mathcal{H}_{2}$ by a linear algorithm $A_{n}:\mathcal{H}_{1}\to\mathcal{H}_{2}$. This is a typical setting in information-based complexity, see, e.g., (EW2008_book1, Section 4.4). Then, the error has the following upper bound for any $g\in\mathcal{H}_{1}$:
@@
In Theorem 2.9, the choice of the cut-off parameter $T$ depends on the smoothness $\alpha$ of the approximand function. However, this information may not be available in practice. Following (KSG2022, Corollary 4.4) and (GKS2023, Remark 3.11), one can replace $\alpha$ by a slowly increasing function $\gamma(n)$ such as $\max(\ln(\ln n),0)$, and still achieve the convergence rate $n^{-\alpha}$, up to a factor $(\gamma(n)\ln n)^{\frac{\alpha}{2}+\frac{1}{2}+\frac{1}{2p}}$.
@@ 3 Spline algorithms
In this section we consider a more refined decomposition of the real line. By partitioning the line into intervals of unit length and having the number of points placed within each interval decrease exponentially fast as one moves away from the origin, we construct an algorithm that attains the optimal rate $n^{-\alpha}$ in $W_{\rho}^{\alpha,q}$ when $q=2$. This construction is similar to (and inspired by) those used in DK2023 and (Karvonen2021, Section 3).
@@ 3.1 Generic results
for all $f\in W^{\alpha,q}(\mathbb{I}_{1})$ and $\nu\in\mathbb{N}$. Let $\mathbb{I}_{k}=[k-1,k]$ for $k\geq 1$ and $\mathbb{I}_{k}=[-k,-k+1]$ for $k\leq-1$. By straightforward translation we obtain algorithms $A_{\nu}^{\mathbb{I},k}\colon W^{\alpha,q}(\mathbb{I}_{k})\to L^{p}(\mathbb{I}_{k})$ such that
@@
which completes the proof. ∎
@@
for all $f\in W^{\alpha,q}(\mathbb{I}_{1})$, where $(x)_{+}=\max\{0,x\}$. See, for example, Section 1.3.11 in Novak1988 or KriegSonnleitner2023; NovakTriebel2006. Because we assume that $1\leq p<q<\infty$, Theorem 3.2 and (3.4) yield the following corollary, which gives the known upper bound on approximation error in $W_{\rho}^{\alpha,q}$ (DK2023, Theorem 3.3).
@@ 3.2 Spline smoothing
where $\widehat{f}$ and $\widehat{\Phi}$ are the Fourier transforms of $f$ and $\Phi$ (see also Theorem 10.12 in Wendland2005). Suppose that $\alpha=\gamma+1/2$ is an integer. Because the Fourier transform of the function $\Phi_{\gamma}$ in (3.6) is proportional to $(c_{1}+c_{2}\omega^{2})^{-(\gamma+1/2)}$ for positive $c_{1}$ and $c_{2}$ (Wendland2005, Theorem 6.13), the space $W_{\gamma}(\mathbb{R})$ in which $K_{\gamma}$ is reproducing on $\mathbb{R}$ has the squared norm
@@
When one has access to the reproducing kernel $K$ of $W(\mathbb{I}_{n})$, constructing the spline smoother is straightforward, though not necessarily computationally convenient. It is a standard result (Wahba1990, Section 1.3) that the minimiser of (3.5) takes the form
@@
Because the Gaussian density function $\rho$ defines a finite measure on $\mathbb{R}$, we have $L^{q}_{\rho}\subseteq L^{2}_{\rho}$ and consequently $W_{\rho}^{\alpha,q}\subseteq W_{\rho}^{\alpha,2}$. The claim then follows from (3.8) and the fact that for points that satisfy (3.9) we have $h_{\nu}\leq C_{h}n^{-1}$ for a positive constant $C_{h}$ independent of $\nu$. ∎
In fact, the smoothness of the Sobolev space that is used to construct the spline smoothers $s_{f,\nu,\lambda_{\nu}}$ does not have to coincide with the smoothness of $f$. It is a consequence of the escape theorem of Narcowich, Ward and Wendland (Narcowich2006, Theorem 4.2) that the estimate (3.8) remains valid even when the Hilbert space $W(\mathbb{I}_{1})$ is norm-equivalent to $W^{\tilde{\alpha},2}(\mathbb{I}_{1})$ for $\tilde{\alpha}\geq\alpha$. From this we obtain the following corollary similar in spirit to Corollary 2.11.

### v3

@@ 1 Introduction
for some $\alpha\in\mathbb{N}$ and $p<q<\infty$. This space consists of functions whose weak derivatives up to order $\alpha$ are in $L_{\rho}^{q}$. Since every element of this space admits a continuous representative, in what follows we always consider the continuous representative so that the function evaluation makes sense.
Function approximation in high dimension is an important task. The sampling recovery problem in Gaussian Sobolev spaces has been recently studied by a number of authors LPE2023 ; DK2023 ; GHRR2023 . In this paper, we study sampling recovery in the one-dimensional Gaussian Sobolev spaces $W_{\rho}^{\alpha,q}$ and construct explicit linear algorithms that attain the optimal or almost optimal rate of worst-case convergence. It has been shown by Kuo, Plaskota, and Wasilkowski (KPW2016, , Theorem 3), and later independently studied by Dũng and Nguyen (DK2023, , Theorem 3.3), that the best possible convergence rate of sampling recovery in $W_{\rho}^{\alpha,q}$ for $p<q$ is of order $n^{-\alpha}$, and we show two alternative methods achieving this rate with explicit computational cost. As one-dimensional algorithms often serve as a foundation for their high-dimensional counterparts, our results motivate further study of such optimal algorithms in a high-dimensional setting.
@@
Numerical integration is closely related to function approximation. Here we mention results related to numerical integration over Gaussian Sobolev spaces for $q=2$. For the deterministic worst-case error, it has been shown that the rate $n^{-\alpha}$ is optimal by Dick et al. (DILP2018, , Theorem 1), meaning that any deterministic linear quadrature cannot achieve a convergence rate faster than $n^{-\alpha}$. We note that their result is shown for so-called Hermite spaces, but the norm equivalence between Hermite spaces and Gaussian Sobolev spaces with $q=2$ is shown in (DILP2018, , Lemma 6) and (KSG2022, , Lemma 2.1). Recently, Dũng and Nguyen (DK2023, , Theorem 2.3) showed that this rate is unchanged for general $q\in(1,\infty)$. The measure being Gaussian, Gauss–Hermite quadrature is a natural choice. However, Kazashi, Goda and one of the present authors KSG2022 have shown that Gauss–Hermite quadrature achieves the deterministic worst-case rate of only $n^{-\alpha/2}$. An upper bound of order $n^{-\alpha/2}$ was essentially obtained already by Mastroianni and Monegato MM1994 and the matching lower bound was proved in (KSG2022, , Theorem 3.2). In contrast, a suitably truncated trapezoidal rule is shown to achieve the optimal rate $n^{-\alpha}$ up to a logarithmic factor in (KSG2022, , Theorem 4.5). The subsequent work of these authors GKS2023 considers the randomized setting, i.e., algorithms are allowed to be random and their quality is measured in the worst-case root-mean-squared error (RMSE). Therein it is proved that no nonlinear adaptive algorithm can converge with a worst-case RMSE rate faster than $n^{-\alpha-1/2}$, and a randomized trapezoidal rule is shown to attain this rate. The strategy used for obtaining upper bounds of trapezoidal rules both in KSG2022 ; GKS2023 is based on introducing an auxiliary periodic function. This strategy is first used in NS2023 , and we adapt this strategy for function approximation in Section 2.
Upon completing this research in the paper, we learned the paper D2025 by Dũng. The methods presented therein achieve the optimal rate of convergence, however they depend on various indices. We emphasize that our methods presented in this paper offer computationally cheap alternatives, especially the trigonometric interpolation. As an additional feature, our methods can be constructed without knowledge of the exact smoothness of the target function. Section 4 concludes this paper by briefly discussing known lower bounds, optimality, and other features of our algorithms.
@@ 2 Trigonometric interpolation
This section introduces our first algorithm based on trigonometric interpolation and analyses its $L^{p}_{\rho}$-error in terms of the number $n$ of function evaluation. The algorithm is denoted by $A^{\dagger}_{n}:W^{\alpha,q}_{\rho}\to L^{p}_{\rho}$.
@@
First we note that $1/q-1/p+\varepsilon<0$. In the following, we use the Sobolev inequality for $F_{\tau}(x)\coloneqq\rho^{1/q-1/p+\varepsilon}(x)g^{(\tau)}(x)$. That is, from the boundedness of $\|F_{\tau}\|_{L^{q}(\mathbb{R})}$ and $\|F_{\tau}^{\prime}\|_{L^{q}(\mathbb{R})}$ we deduce the boundedness of $\|F_{\tau}\|_{L^{\infty}(\mathbb{R})}$. Let $H_{\ell}(x)$ be the $\ell$-th degree probabilist’s Hermite polynomial,
@@
First notice from the definition of Hermite polynomials that
@@
where in the last line, again we used $\|f^{(\tau-\ell+1)}(x)\|_{L^{q}_{\rho}}\leq\|f\|_{W^{\alpha,q}_{\rho}}<\infty$. Thus we have $\|g\|_{\textup{decay}}<\infty$. ∎∎
@@
The proof follows the strategy of (KSG2022, , Lemma 4.1), where the case with $p=1$ and $q=2$ was proved. Recall that for $\tau,\tau^{\prime}\in\mathbb{N}_{0}$ scaled Bernoulli polynomials satisfy
@@
Next, we note that $G^{(\alpha)}$ and $g^{(\alpha)}$ only differ by the constant $\frac{1}{2T}\int_{-T}^{T}g^{(\alpha)}(y)\,\mathrm{d}y$ on [T,T][-T,T]. Since $f^{(\alpha)}\in L^{q}_{\rho}$, we know $g^{(\alpha)}$ is in $L^{q}([-T,T])$. Due to matching boundary values $G^{(\alpha-1)}(-T)=G^{(\alpha-1)}(T)$ and absolute continuity of $G^{(\alpha-1)}$ on $(-T-\delta,T+\delta)$, we have $\|G^{(\alpha)}\|_{L^{p}(\mathbb{T}([-T,T]))}<\infty$. ∎∎
@@
where we used the facts that $H_{\ell}(\cdot/\sqrt{p})\rho^{(q-p-\varepsilon)/p}$ is bounded on the real line and that $\rho(x)$ is decreasing in $\lvert x\rvert$. ∎∎
@@
The claim then follows from Lemma 2.4 with a=Ta=-T, $b=T$, and $\varepsilon=0$. ∎∎
@@
The result by Temlyakov (T2018_book, , Theorem 2.4.4), where the interpolation operator is denoted by $I_{m}$, applies to our setting. Therein, the function space considered is defined via convolution kernels, but the norm equivalence to periodic Sobolev spaces
@@
is shown in (U2007, , Theorem 2.7). Hence, by letting $S_{T}(x)\coloneqq Tx/\pi-T$, (T2018_book, , Theorem 2.4.4) leads to
@@
Thus we have proved the claim. ∎∎
@@
where in the penultimate line, we used $|\frac{B^{[-T,T]}_{\tau}(x)}{\tau!}|\leq\frac{(2T)^{\tau-1}}{2}$ for $x\in[-T,T]$; see (NS2023, , Equation (6)) or L1940 .
@@
where this explicit constant here can be found, e.g., in (SN2016, , Theorem 2.1), or we refer to EZ1966 as an earlier result.
@@
Hence we proved the claim.∎∎
@@
This proves the claim.∎∎
@@
$\|f-A^{\dagger}_{n}(f)\|_{L_{\rho}^{p}}\leq C_{p,q,\alpha,\varepsilon}\|f\|_{W^{\alpha,q}_{\rho}}(\ln n)^{\frac{\alpha}{2}+\frac{1}{2}+\frac{1}{2p}}n^{-\alpha}.\qed$
@@
This proof strategy of introducing the auxiliary function $G$ can be applied to other algorithms or problems. Let us elaborate this part here. Consider an abstract problem of approximating a linear operator or a linear functional $S:\mathcal{H}_{1}\to\mathcal{H}_{2}$ by a linear algorithm $A_{n}:\mathcal{H}_{1}\to\mathcal{H}_{2}$. This is a typical setting in information-based complexity, see, e.g., (EW2008_book1, , Section 4.4). Then, the error has the following upper bound for any $g\in\mathcal{H}_{1}$:
@@
In Theorem 2.9, the choice of the cut-off parameter $T$ depends on the smoothness $\alpha$ of the approximand. However, this information may not be available in practice. Following (KSG2022, , Corollary 4.4) and (GKS2023, , Remark 3.11), one can replace $\alpha$ by a slowly increasing function $\gamma(n)$ such as $\max(\ln(\ln n),0)$, and still achieve the convergence rate $n^{-\alpha}$, up to a factor $(\gamma(n)\ln n)^{\frac{\alpha}{2}+\frac{1}{2}+\frac{1}{2p}}$.
@@ 3 Spline algorithms
In this section we consider a more refined decomposition of the real line. By partitioning the line into intervals of unit length and having the number of points placed within each interval decrease exponentially fast as one moves away from the origin, we construct an algorithm that attains the optimal rate $n^{-\alpha}$ in $W_{\rho}^{\alpha,q}$ when $q=2$. Naive version of this construction based on solving a linear system [see (3.7)] requires $\mathcal{O}(n^{3})$ computational cost, and is similar to (and inspired by) those used in DK2023 and (Karvonen2021, , Section 3). Faster computation is possible in some special cases that we do not discuss here.
@@ 3.1 Generic results
for all $f\in W^{\alpha,q}(\mathbb{I}_{1})$ and $\nu\in\mathbb{N}$. Let $\mathbb{I}_{k}=[k-1,k]$ for $k\in\mathbb{Z}$. By straightforward translation we obtain algorithms $A_{\nu}^{\mathbb{I},k}\colon W^{\alpha,q}(\mathbb{I}_{k})\to L^{p}(\mathbb{I}_{k})$ such that
@@
which completes the proof. ∎∎
@@
for all $f\in W^{\alpha,q}(\mathbb{I}_{1})$, where $(x)_{+}=\max\{0,x\}$. See, for example, Section 1.3.11 in Novak1988 or KriegSonnleitner2023 ; NovakTriebel2006 . Because we assume that $1\leq p<q<\infty$, Theorem 3.2 and (3.4) yield the following corollary, which gives the known upper bound on approximation error in $W_{\rho}^{\alpha,q}$ (DK2023, , Theorem 3.3).
@@ 3.2 Spline smoothing
where $\widehat{f}$ and $\widehat{\Phi}$ are the Fourier transforms of $f$ and $\Phi$ (see also Theorem 10.12 in Wendland2005 ). Suppose that $\alpha=\gamma+1/2$ is an integer. Because the Fourier transform of the function $\Phi_{\gamma}$ in (3.6) is proportional to $(c_{1}+c_{2}\omega^{2})^{-(\gamma+1/2)}$ for positive $c_{1}$ and $c_{2}$ (Wendland2005, , Theorem 6.13), the space $W_{\gamma}(\mathbb{R})$ in which $K_{\gamma}$ is reproducing on $\mathbb{R}$ has the squared norm
@@
When one has access to the reproducing kernel $K$ of $W(\mathbb{I}_{n})$, constructing the spline smoother is straightforward, though not necessarily computationally convenient. It is a standard result (Wahba1990, , Section 1.3) that the minimiser of (3.5) takes the form
@@
Because the Gaussian density function $\rho$ defines a finite measure on $\mathbb{R}$, we have $L^{q}_{\rho}\subseteq L^{2}_{\rho}$ and consequently $W_{\rho}^{\alpha,q}\subseteq W_{\rho}^{\alpha,2}$. The claim then follows from (3.8) and the fact that for points that satisfy (3.9) we have $h_{\nu}\leq C_{h}n^{-1}$ for a positive constant $C_{h}$ independent of $\nu$. ∎∎
In fact, the smoothness of the Sobolev space that is used to construct the spline smoothers $s_{f,\nu,\lambda_{\nu}}$ does not have to coincide with the smoothness of $f$. It is a consequence of the escape theorem of Narcowich, Ward and Wendland (Narcowich2006, , Theorem 4.2) that the estimate (3.8) remains valid even when the Hilbert space $W(\mathbb{I}_{1})$ is norm-equivalent to $W^{\tilde{\alpha},2}(\mathbb{I}_{1})$ for $\tilde{\alpha}\geq\alpha$. From this we obtain the following corollary similar in spirit to Corollary 2.11.
@@
4 Optimality
We conclude this paper by discussing optimality, i.e., general lower bounds for algorithms which are based on function values. For $p<q$, the optimal rate is derived by Kuo, Plaskota, and Wasilkowski (KPW2016, , Theorem 3). They consider a more general class of weights, which is applicable to ours. Their result states that the error of any algorithm $A_{n}$ that only uses $n$ function evaluation has the lower bound
$\sup_{\begin{subarray}{c}f\in W^{\alpha,q}_{\rho}\\ \|f\|_{W^{\alpha,q}_{\rho}}\leq 1\end{subarray}}\|f-A_{n}(f)\|_{L_{\rho}^{p}}\geq C_{p,q,\alpha}n^{-\alpha},$
for $1\leq p<q$. In this sense, our methods achieving error convergence $n^{-\alpha}$ are optimal. In the same paper, they propose a piecewise Lagrange approximation that achieves the optimal rate (KPW2016, , Theorem 2).
Even though our algorithms do not cover the extreme case $p=q$, this has been recently studied by Dũng (D2025, , Theorem 3.4). In this case the optimal rate degenerates to $n^{-\alpha/2}$. Dũng D2025 also covers extreme cases, such as $q=1$ or $q=\infty$, deriving the convergence rate $n^{-\alpha/2+(1-1/p)/2}$ for $q=1$ and $n^{-\alpha/2+1/(2p)}$ for $q=\infty$. These rates can be achieved by his B-spline algorithms (D2025, , Theorem 2.2 and 2.4).
In this paper, we focused on implementable and efficient algorithms achieving the optimal approximation rate. In practice, we often do not know the smoothness $\alpha$ of the target function and we have access to only function values. We explored the case when $\alpha$ is not known exactly in Corollaries 2.11 and 3.6. It is an interesting question of whether optimal algorithms can be constructed for the extreme cases of $(p,q)$ without knowledge of $\alpha$. We leave this to future study.